OpenStudy (anonymous):

Solve sinθcosθ-cosθ=0 for 0≤θ<2π

OpenStudy (anonymous):

I have done the following: cosθ(sinθ-1)=0 Using Zero-Product Property either cosθ=0 which gives θ=π/2 or sinθ-1=0 ==> sinθ=1 which gives θ= π/2 Are there other possible values of the angle θ within the given range?????

OpenStudy (anonymous):

\[ \cos(\theta) = 0 \rightarrow \frac{\pi}{2} \text{ or } \frac{3\pi}{2} \]